OpenStudy (loser66):

An odd polynomial is an odd degree polynomial but not vice versa. How about a completely odd degree polynomial? I mean how about a polynomial whose terms are odd degree like x^5 + x^3 + x? Is it an odd polynomial for sure? Please, explain me

OpenStudy (loser66):

@mathmale

OpenStudy (mathmale):

Hello, Winner66! If all the exponents of x in the polynomial in question are odd, then I think you can safely label it as an odd polynomial. Are you familiar with the following characteristic of an odd poly.? f(-x) = -f(x). In what context did you find this problem?

OpenStudy (loser66):

Yes, I am familiar with odd poly definition My problem come from my test which I failed when confusing between odd poly and odd degree poly. My problem on test was: Let U \(\subset P_4\) be the subspace of all polynomials x \(\in P_4\) such that x'(0)=0. Let W\(\subset P_4\) be the subspace of all odd polynomials. a) What is U+W? justify your answer b)What is U\(\bigcap\)W? justify your answer c) Write the polynomial x = 1+ t=t^2 +t^3 as a sum of u + w where u \(\in U \) and w \(\in W\) in 2 different ways. I failed. hihihihi...

OpenStudy (loser66):

for c * 1+ t+t^2 + t^3

OpenStudy (mathmale):

An odd polynomial is one whose every power of x is an ODD integer. An odd-order polynomial is one whose FIRST TERM involves an ODD power of x.

OpenStudy (loser66):

Got it, thank you so much.

OpenStudy (mathmale):

I'm really happy to have been able to help with this. All the best to you.

OpenStudy (loser66):

Thank you :)